OpenStudy (anonymous):

An acute angle θ is in a right triangle with sin θ =2/3 . What is the value of cot θ? I found that 2 is the opposite angle and 3 is the hypotenuse. adj/opp = cot θ, so adj/2 = cot θ I tried using the Pythagorean Theorem, but I can get an answer choice! Please help! Thanks!

jimthompson5910 (jim_thompson5910):

The adjacent side can be found by the Pythagorean Theorem a^2 + b^2 = c^2 x^2 + 2^2 = 3^2 x^2 + 4 = 9 x^2 = 9-4 x^2 = 5 x = sqrt(5) So the adjacent side is sqrt(5) units which means cot(theta) = adj/opp cot(theta) = sqrt(5)/2

jimthompson5910 (jim_thompson5910):

Here's a better look at the answer \[\Large \cot(\theta) = \frac{\sqrt{5}}{2}\]

OpenStudy (anonymous):

Huh! So I did do it right! I must have done the math wrong! Thanks for the clarification! :)

jimthompson5910 (jim_thompson5910):

yw